OpenStudy (anonymous):

Expand and evaluate the power. 28 A. 16 B. 256 C. 512 D. 1024

OpenStudy (anonymous):

2 to the 8th power

OpenStudy (anonymous):

2^8 = 2*2*2*2*2*2*2*2 = 256